Write an equation of the line that passes through (8,-5) and is perpendicular to the graph of the line that passes through points (3,5) and (4,9).
Answer by lwsshak3(11628)

Write an equation of the line that passes through (8,-5) and is perpendicular to the graph of the line that passes through points (3,5) and (4,9).
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
slope(m)=∆y/∆x=(y1-y2)/(x1-x2)
For the line that passes through points (3,5) and (4,9).
m=(9-5)/(4-3)=4/1
..
For the line that passes through (8,-5), perpendicular to the first line
m=-1/4 (negative reciprocal of slope of first line)
equation: y=(-1/4)x+b
solve for b using given coordinates(8,-5) on the line
-5=(-1/4)*8+b
-5=-2+b
b=-3
equation: y=-x/4-3